OpenStudy (anonymous):

solve each system of equations by substitution c+2d=12 c+d =2

OpenStudy (anonymous):

alright, look at the second equation 1) let's get d by itself

OpenStudy (anonymous):

so d=2-c?

OpenStudy (anonymous):

c=2-d and then replace c in the first equation 2-d+2d=12 => d=10 then replacing it in c c=-8

OpenStudy (anonymous):

good now in first equation , put that inplace of d

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

so d=2-c now substitute into first equation c+2d=12 c+2(2-c)=12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got c =8 im way off

OpenStudy (anonymous):

c+4-2c=12 -c+4=12 -c=12-4 -c=8 c=-8

OpenStudy (anonymous):

ermagerwd im just getting more confused

OpenStudy (anonymous):

its just confusing me more

OpenStudy (anonymous):

A: c+2d=12 |>> in B c+d=2 = c=2-d substitute in A that c=2-d we will get B: c+d=2 | -we will get in A c+2d=12 ;(c=2-d) then c in A turn to 2-d means 2-d+2d=12 = 2+d=12 = d=12-2 = d=10 <get one substitute in B c+d=2 ;(d=10) then c+10=2 = c=2-10 = c=-8 d=10 c =-8 i hope you not confuse now

OpenStudy (anonymous):

haha yeah right im definatly confused

OpenStudy (anonymous):

can u read my answer. it is simple

OpenStudy (anonymous):

yeah but its confusing me all of this is i fluttering hate math it makes no fluttering sense to me

OpenStudy (anonymous):

it seems that u did not but ur effort on it.